UT 27 Aug 2020 - Night #6 of Restart

OSA: S. Allanson

ISp: C. Bolyard

Mtn Manager: L. Durham

ISA: O. Kuhn (Tucson)

Special Guest Stars: B. Rothberg (Tucson), J. Power (Tucson)

Summary:

We were open for most of the night, but got a late start due to clouds and then closed briefly due to smoke.

LUCI2:
  • Rotator zeropoint: We started with LUCI2 and Jenny did the 3 MOS alignments which used the 1" slit + AFC holes and a PA such that two stars would fall in the slit. The rotational offset needed to align the mask was about 0.1 deg for two fields and ~0 for the other, however for that one the confirmatory thru-slit image did not have the two stars as well aligned along the slit as for the other fields.
  • Rotator center on M52: Next we obtained data to measure the rotator center on the detector
  • Science like imaging on M52
  • Non-sidereal imaging observation
LBCR:
  • NS imaging with guiding and a dither
  • Science like imaging on M76 to complete the set of images through all filters
  • Rotational trailing test on a field just after transit with rich logging of rotator trajectories. Trailing was not seen on the first image (or was not very easy to see), but it was obvious on the subsequent image and the arcs decreased in length as the object moved further from zenith. This is as expected for a timing/RA error. The observing night when this was noted in January seemed to have more of a mix of trailed and not trailed images.
  • Twilight sky flats at R-Bessel & F972N20. Last night and tonight, we obtained good flats at i, r, R, F972N20 and Y. We still need z, V and I.

PROBLEMS

IT 8171: MODS1 Blue grating timeout error. I encountered this during coldStart on Aug 20, but at that time a reset of the microLynx controller (as in IT 7935) resolved the issue. We used MODS1 almost all night last night and cycled through all 3 position at least twice when observing standards in all modes and, before that, when running focus sequences. Tonight, a reset of the individual controller is not doing the trick. MODS1 is offline until this is fixed.

LBCR: the spot seen on last nights images of M76, but not on the flats, was not seen tonight.

LBCR flats: I noticed that the table of flats is not auto-populating. This could be a consequence of the archive issues. I'll look into it a bit and ask Yang.

Night Log:

UT 00:21 - still troubleshooting AGW-1

UT 02 - sieveSnaps on both MODS. MODS1 Blue Grating problem (see above; IT 8171). But MODS2 Blue images are not showing evidence of sequencer issues as they were (7 in a row) yesterday (IT 6386).

UT 06 - Steve is opening - the clouds have finally cleared out. The first task is to check the rotator zero point. Jenny is online to do this alignment as she designed the scripts/LMS files to do the MOS alignment using unboxed reference stars and the 1" long slit. Two fields are available at this time of the night: Stone O and Stone P. Going to Stone O first.

UT 06:07 Slew to Stone O field. Pointing and collimation.
  • No star appeared in the acquisition image (wrong starting IE/CA values from old night report, ok after noticed).
  • Collimation - the "bite" is back (or still there)
  • Check FS alignment using AFC spots above/below slit - only a few pixels off.
UT ~06:20 Run script: Alignment images: sky,obj,slit = 66,67,68.
  • error - auto-mode mask alignment failed, NULL exception. requested plane not available for LMS target. Jenny is working on it. She found the problem - a MOS alignment with an actual MOS mask needs to be done in the RTD before a MOS alignment with a LS+AFC holes is done
UT 06:30 Steve is closing because the 25-mic particulate counts reached the limit.

UT 07:02 Steve is reopening as the particle counts dropped.

UT 07:16 Trying again with the alignment:
  • sky,obj,slit = 71,72,73. The mask alignment on detector ok. and then the actual to predicted positions of the reference stars were all different by 10" which was the offset to sky. Reversing the assignment of sky and object files helped.
  • (x-shift,y-shift,angle = 0.09736", -3.10276", 0.01454deg) error RA/DEC 0.06740"/0.15193"
  • angle offset is 0.01454 deg. The confirmatory image had a slight tilt(misalignment) of the stars in the slit
UT 07:27 Now trying the same on Stone P
  • sky,obj,slit = 75,76,77. This time, reversing the object, sky was not needed.
  • (x-shift,y-shift,angle) = -0.22575, -3.48477, 0.12771 deg, error 0.09710"/09592"
  • angle offset is 0.12771 deg. Confirmatory looks better.
UT 07:50 Stone A
  • sky,obj,slit = 79,80,81. This time, reversing object,sky not needed. Only time reversing needed was when images were manually loaded - related?
  • (x-shift,y-shift,angle)= -0.02005, -3.74295, 0.11710 deg, error 0.0586"
  • angle offset is 0.11710 deg, similar to Stone P. Confirmatory looks better.
Results are that the rotator zeropoint seems to be about ~0.1deg off.

UT 08:07 Rotator Center on Detector
  • M52 active 83,84 (object,sky)
  • track 0 85
  • track 180 86
  • track 270 87
  • track 90 88
UT 08:25 Imaging script on M52
  • Paused after 1st exposure to realign FS (this is image 90).
  • J 89,91-99
  • K 100-102
UT 08:54 Non-sidereal imaging on Comet 29/P Schlassmann-Wachmann 1
  • K 103-111
  • H 112-116
  • J 117-124
UT 09:30 Steve is reconfiguring for LBCR. He had to retract SX M3 in order to do the dynamic balance.

UT 09:55 Slewed to the field of a PHA for a test of NS guiding.
  • Collimating on the co-pointing field. There is a lot of coma. WRS did not correct it and the first FPIA correction sent 3000 nm Z8. This was like last night, and needs to be looked into. Eventually converged in 5 FPIA iterations.
  • Pointing correction using ptautoadjust (display scaling is awful)
  • Running NS OB on 2018 UQ1. The OB will take 3 180-sec exposures and each of 2 dithers. r-SLOAN. On 101006, at 951, 4103. 26K counts... 180s was plenty, 120s would have sufficed.
  • It is guiding, but the object and the PSF are elongated.
  • Images: lbcr.20200827.hhmmss.fits, hhmmss=
    • 101006, 101344, 101722 at 0,0
    • 101722, 102107, 102444 at a small offset
UT 10:37 Slewing to M76 for more guiding+dithering sidereal tests (yesterday @ rizR, today at VIFY)
  • dofpia, /redonly
  • ptautoadjust on M76
  • 03:51 UT: I started an OB to cycle through the riz filters
  • blinking the z images from last night and tonight, it can be seen that the spot was only on last night's image. Remains to examine the other filters.
  • 04:02 UT: I started an OB to cycle through VRIFY filters.
UT 11:21 Rotator trailing test to collect log data on-sky
  • Focus/Collimate, shift-S to stop because the coeffs are small but est seeing is too and the coeffs are just above the criteria
  • ptautoadjust - some IRAF-ds9 problems req restart of ds9 and iraf. Ultimately, offsets sent were 4, 14.
  • running an OB with large dithers and for coordinates of a field that has just transited and is at 86 deg elevation.
    • the PSFs are elongated and there may be a sign of rotational trailing but it is not too easy to tell at first glance.
    • on 114527, there is clearly rotational trailing; and
    • on 114910, the PA is definitely changing during the observation, and there is rotational trailing.
    • on 115255, el 84:45, trailing
    • diminishing trailing as we move farther from the zenith
UT 12:10 Slewing to field for skyflats.
  • got a good set of R-Bessel flats at PA=0 and PA=180
  • onto F972N20 --- a set of ~5-7 good flats here, mostly at PA=180.
UT 12:41 Closing up.
  • I put both MODS to sleep
  • Taking LUCI2 imaging flats:
    • J 121-125, 126-130
    • H 131-135,136-140
    • K 141-145,146-150
  • Taking a series of LBCR biases.
  • LUCI is in an end-of-night state: N3.75, no mask (last cals)
  • Turning off LBCR. I left HK Red on, but HK Blue is OFF

Proposed Plan

  • LUCI-2
    • Rotator Center with N30 fieldstop (0.5 hr) - high
    • Rotator ZP check using Longslit method (0.5hr) - high
    • Transform verification (1hr) - medium (could also wait until LUCI-1 is on and take thorough data set, but 2x work for D. Miller)
    • Imaging Check
      • short OB (M52 three filters - 0.5-1hr) - sidereal - high
      • J1601 -K only 15-20min - sidereal - only if really good seeing.
      • NGC 7027 - includes all filters -- 1.5 hrs
      • NGC 40 - includes all filters - 1.5 hrs
    • Imaging Check - non-sidereal
      • 29P Comet (like MODS) - high - 0.5hr (JHK filters)
  • LBC-Red checkout
    • NS OB (targets TBD)
    • Mixed Mode with MODS-1
    • Standard fields (weather permitting)
    • Spectrophotometric Stds (weather permitting)
    • Skyflats (skies permitting)
  • AGW-1 hotspot centering (no WFS)
-- BarryRothberg - 26 Aug 2020
Topic revision: r10 - 29 Aug 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback